A particle executes SHM of amplitude 5m when the particle is 3m from its mean position, its
acceleration is found to be 48m/s2
. Find (i) velocity (ii) time period (iii) Maximum velocity
1
Expert's answer
2021-08-10T17:54:56-0400

"v=w\sqrt{a^2-x^2}"

"v=w\sqrt{5^2-3^2}=w\sqrt{25-9}=w\sqrt16=4w"

"a=w^2x=w^2\times3"

"48=3w^2\\w^2=16"

"w=4rad\/sec"

"T=\frac{2\pi}{4}=1.57sec"

V=aw

"v_{max}=5\times4=20"